### v1

# Recov-Vision: Linking Street View Imagery and Vision-Language Models for Post-Disaster Recovery

Abstract
Building-level occupancy after disasters is vital for triage, inspections, utility re‑energization, and equitable resource allocation. Overhead imagery provides rapid coverage but often misses facade and access cues that determine habitability, while street‑view imagery captures those details but is sparse and difficult to align with parcels. We present FacadeTrack, a street‑level, language‑guided framework that links panoramic video to parcels, rectifies views to facades, and elicits interpretable attributes (for example, entry blockage, temporary coverings, localized debris) that drive two decision strategies: a transparent one‑stage rule and a two‑stage design that separates perception from conservative reasoning. Evaluated across two post–Hurricane Helene surveys, the two‑stage approach achieves precision 0.927, recall 0.781, and F‑1 score 0.848, compared with the one‑stage baseline at precision 0.943, recall 0.728, and F‑1 score 0.822. Beyond accuracy, intermediate attributes and spatial diagnostics reveal where and why residual errors occur, enabling targeted quality control. The pipeline provides auditable, scalable occupancy assessments suitable for integration into geospatial and emergency‑management workflows.
I Introduction
I-A Motivation & problem statement
Rapid, building-level occupancy status after a disaster is a time-critical input for triage, utility restoration, inspections, temporary shelter planning, and equitable resource allocation. Overhead (nadir) imagery arrives quickly and covers large areas, but its roof-dominant perspective often overlooks the very cues that determine whether people can safely inhabit a building–open or blocked entries, temporary coverings, localized debris, and ad-hoc repairs. Even state-of-the-art post-disaster overhead methods struggle when habitability hinges on facade and access conditions rather than roof damage alone [1, 2, 3]. Street-based visual surveys and door-to-door assessments supply those missing signals, yet they are inherently sparse and viewpoint-dependent, constrained to the road network and acquisition schedules, and their alignment with parcel inventories is noisy in practice [4, 5, 6, 7]. As a result, analysts often face a slow, manual reconciliation of overhead and ground evidence that is difficult to reproduce or audit at scale, and that typically occurs only once immediately after impact rather than being repeated to track recovery over time.
Recent progress in vision-language models (VLMs) could provide a mechanism for turning heterogeneous visual evidence into operational decisions without hand-crafted taxonomies for every new event. By combining open-vocabulary recognition, instruction following, and cross-modal reasoning, VLMs can be prompted to connect small visual cues with higher-level statements about building condition and use [8]. Early domain examples point in this direction—for instance, explanation-driven change reasoning on paired remote-sensing images [9], multi-hazard benchmarks for multimodal analysis [10], and large-model-assisted enrichment of facade attributes [11]. These studies establish feasibility and hint at auditability through natural-language rationales, but they stop short of a reproducible, street-level pipeline centered on post-disaster occupancy inference.
A key limitation of past works is stitching together complementary evidence across views, turning ambiguous signals into explicit intermediate attributes, and proving reliability against specialized baselines. Overhead and street-level perspectives capture different failure modes–roof versus entry and facade–and simple late fusion rarely suffices when coverage is uneven or views are occluded [5]. Small, indirect indicators such as vehicle mix, tarps behind foliage, or shadows around doorways need to be promoted into structured, inspectable attributes so that an assessor can see how the model moved from evidence to judgment [11, 8]. And because domain-tuned aerial and street-view methods already achieve high accuracy–often in the face of severe class imbalance–claims of progress require statistical comparisons, spatial localization of errors, and analysis of failure cases tied to viewpoint and occlusion [3, 1, 2, 4]. These gaps become more pressing each year as large-scale flood and storm disasters affect growing coastal and urban populations, and as agencies need scalable, auditable assessments to guide and defend resource-allocation decisions [12].
Agencies need parcel-level occupancy within days to triage inspections, re-energize utilities, plan temporary shelter, and allocate resources equitably. Yet nadir imagery systematically misses facade and access cues—blocked entries, emergency placards, tarps, exterior mud—that determine habitability, while ad-hoc reconciliation of ground photos is slow and hard to audit. The study’s rationale is to close this gap with a reproducible, street-level, language-guided pipeline that makes evidence-to-judgment traceable: it links 5.6K 360° video to parcels, estimates heading from short GPS windows, and uses a simple yaw formulation to generate facade-centered rectified views. It then elicits nine interpretable visual attributes and applies two complementary decision strategies—a transparent one-stage rule with vehicle adjustment and a two-stage design that separates perception from conservative reasoning—so operating policies can be tuned without retraining. Together, these choices address the core needs of auditability, scalability, and defensibility for post-disaster occupancy decisions.
I-B Research objectives and significance
This work seeks to develop a street-level, language-guided framework for post-disaster building occupancy assessment that combines street-level imaging, automated facade rectification, and vision-language reasoning to produce parcel-level occupancy labels and recovery trajectories. Our objective is to develop a reproducible street-level pipeline that combines facade rectification, VLM attribute extraction, and conservative reasoning to monitor parcel-level occupancy change across repeat visits. This objective set emphasizes operational reproducibility, interpretable intermediate evidence, and statistically grounded temporal and spatial evaluation for recovery monitoring.
[…]
I-C Contributions
Despite major progress in overhead change detection, habitability remains driven by facade and access conditions that nadir views routinely miss—blocked entries, emergency placards, exterior mud, ad-hoc repairs—forcing analysts into slow, ad-hoc reconciliation of roof-dominant evidence with scattered ground photos that are difficult to reproduce or audit. Street-view imagery contains the missing cues but is sparse, viewpoint-dependent, and noisy to align with parcel inventories in practice, while recent VLM efforts demonstrate feasibility without delivering a reproducible, street-level pipeline for post-disaster occupancy inference. What remains unsolved is threefold: cross-view integration when coverage is uneven or occluded, promoting weak, ambiguous signals (such as vehicles behind foliage or mud lines) into structured, inspectable attributes, and rigorous validation against specialized baselines with statistical tests, spatial localization of residual error, and failure-mode analysis under class imbalance. These persistent knowledge gaps motivate a different starting point.
Operationally, agencies need parcel-level occupancy within days to triage inspections and re-energize utilities, but current workflows rely on overhead-only triage or manual, non-auditable fusion of ground evidence, with no standard way to trace evidence to attributes to decision, tune risk-aware thresholds without retraining, report temporal change (recovered/deteriorated) credibly, or direct QA/QC using spatial error diagnostics. In short, practice lacks an end-to-end, repeatable capture-to-label chain—from panoramic video and GPS linkage through facade rectification to interpretable attributes and defensible decisions—that can scale while preserving auditability and privacy. Establishing this tooling and evaluation scaffolding is the point of departure for the present study: to replace ad-hoc reconciliation with a modular, language-guided, street-level pipeline explicitly designed for measurement fidelity, policy tuning, and targeted human oversight.
[…]
I-D Paper organization
Section II reviews (i) vision-language foundations, (ii) aerial/satellite imagery-based damage assessment, and (iii) street view-related pipelines. Section III introduces data collection, frame-parcel linkage, panoramic rectification, attribute labeling, and the two decision strategies. Section IV reports classification performance, representative examples, two-visit occupancy change, and spatial error patterns. Section V discusses deployment implications, limitations, and future directions. Section VI concludes with key contributions and implications for rapid, auditable occupancy monitoring.
II Related work
II-A Aerial and satellite imagery for post-disaster damage
Deep learning with overhead imagery has enabled rapid, wide-area mapping of building damage through semantic segmentation and change detection. Synthesizing this progress, [13] survey architectures, datasets, and open challenges for remote-sensing-based damage detection. Early pipelines used convolutional models on open aerial data to map impacts [14], while operational analyses documented how nadir views are confounded by cloud, canopy, and shadows, which obscure key cues and complicate severity estimation [15]. Alongside dataset curation, typology-driven formulations argue for more actionable, interpretable categories aligned with decision needs rather than coarse severity [16].
Modeling advances improved temporal fusion and representation capacity. Hierarchical transformers scale to large areas while separating fine-grained categories [2]. For example, Siamixformer introduces a fully transformer-based Siamese architecture for bi-temporal change reasoning [17] and BDANet combines multiscale features with cross-directional attention to capture both local and global cues [1]. UAV and oblique-aerial studies extend beyond satellites to quantify post-event damage with tailored models and views [18, 19, 20].
Overall, overhead approaches deliver speed, coverage, and mature change-detection tool chains, and remain indispensable for regional triage. However, they struggle when habitability hinges on entries and facades rather than roof conditions [21], and performance can degrade under severe class imbalance or subtle states [22, 15]. These limitations motivate complementing the nadir evidence with street-level cues and adopting reasoning pipelines that can articulate functional implications at the building scale.
II-B Street-view imagery for urban condition and post-event assessment
Street-view imagery (SVI) provides facade-level detail that reflects access and near-term functionality. A panoramic SVI survey documents broad applications in data-driven urban research and the methodological considerations that come with roadside acquisition [6], while open-source pipelines have been introduced to streamline ingest, filtering, and sampling for research and operations [7]. In post-disaster settings, bi-temporal SVI supports hyperlocal change assessment at the building scale [4]; multi-modal fusion with structured attributes improves granularity and predictive performance for damage outcomes [5].
Coverage and viewpoint variability motivate multi-view designs and fusion with other sensors. Leveraging multiple SVI vantage points improves robustness under occlusions and oblique angles [23], and fusing SVI with remote sensing supports broader risk and exposure modeling in flood contexts [24, 25]. These advances show how ground imagery complements overhead views with human-interpretable evidence, even as road-network constraints and capture schedules limit universal availability.
II-C Vision-language and multimodal foundations for the built environment and disasters
Vision-language models offer natural-language interfaces to visual evidence, enabling open-vocabulary recognition and cross-modal reasoning with limited task-specific training. [8] survey emerging applications in remote sensing; recent work tailors VLMs to disaster analysis and built-environment tasks, enabling models to capture context-specific features such as flood extent, building accessibility, and structural condition that are essential for recovery monitoring. ChangeChat demonstrates explanation-driven change reasoning on paired remote-sensing images [9], and DisasterM3 provides a multi-hazard, multi-sensor benchmark to evaluate multimodal models under diverse conditions [10]. In street-level contexts, language-conditioned estimation of lowest-floor elevation illustrates how VLMs can extract actionable attributes from facade imagery [26], and large-model-assisted enrichment frameworks expand architectural attributes at scale using SVI [11]. Complementary efforts on knowledge fusion and data generation aim to structure damage information and mitigate imbalance by combining imagery with expert priors within VLM-driven pipelines [27].
Taken together, this survey of past studies suggests that language-guided reasoning can connect small, ambiguous cues to decision-ready statements while exposing intermediate rationales for auditing. Prompt sensitivity, small-object perception limits, and the need for domain-native resources remain practical hurdles, but as coastal and urban risks intensify [12], the combination of overhead imaging, SVI, and VLMs offers a promising basis for auditable, building-level occupancy inference. Our work builds on these strengths while addressing the remaining gaps around cross-view integration, transparent attribute-to-decision tracing, and rigorous evaluation against specialized baselines.
III Methodology
This study proposes a pipeline that combines Street View Imagery (SVI) with Vision Language Models (VLMs) to infer building occupancy status in post-disaster scenarios. The method integrates geospatial data collection and processing, computer vision techniques, and natural language reasoning to assess whether residential buildings are occupied following extreme disaster events. The approach consists of four key components: (1) collection and preprocessing of panoramic street-view video data with GPS tracking to generate building-specific facade views, (2) development of a vision-language model prompting strategies that extract visual damage indicators and infer occupancy status, (3) temporal inference modeling to track occupancy changes over time, and (4) comprehensive evaluation against ground truth data. This multi-modal framework enables scalable, objective assessment of post-disaster building conditions using street-level imagery and state-of-the-art language models.
III-A Data collection
As a test case, we collected imagery from two field visits to areas impacted by Hurricane Helene. Helene made landfall as a Category 4 hurricane in the Big Bend region of Florida on September 26, 2024, before traveling inland and stalling over the North Carolina/Tennessee border on September 27–28. In total, Helene is estimated to have generated more than \$78 billion in economic losses [28] and at least 250 fatalities. Of these, 95 were directly attributable to freshwater flooding (which includes landslides and debris flows) in North Carolina, South Carolina, and Tennessee, making it the deadliest hurricane to make landfall in the continental United States since Hurricane Katrina (2005) [29]. Notably, Helene’s heavy precipitation (i.e., 20–30 inches, corresponding to a 1000-year rainfall event in western North Carolina) over a region characterized by steep topography generated widespread flooding, damaging tens of thousands of homes and cutting off entire communities from access to critical facilities such as water, sewer, electricity, telecommunications, and healthcare [30]. The southern Appalachian Mountains exhibit particularly low flood insurance penetration (i.e., only 1.5% of buildings had coverage [31]), raising questions about how quickly the region will recover and what household recovery will look like.
Our study focused on several flooded neighborhoods in Swannanoa and Black Mountain in Buncombe County, North Carolina. (Swannanoa is a census-designated place located between the City of Asheville and the Town of Black Mountain.) Both communities are located along the Swannanoa River, which is a tributary of the French Broad River. The key survey areas are shown in Figure 1. Prior post-disaster surveys (e.g., Red Cross assessments and county emergency management reports [32]) identified these neighborhoods as among the most heavily affected communities in the region. Our field reconnaissance in December 2024 confirmed the high density of visible damage, making them appropriate testbeds for property-level monitoring. The damaged areas comprised a mix of suburban and semi-rural parcels and had relatively accessible road networks, enabling repeated street-level imaging to support temporal analysis of recovery. In addition to the initial reconnaissance visit conducted in December 2024, we collected street-level data across two field campaigns in March 2025 and June 2025.
Figure 1: Study area in broader Asheville area, North Carolina, showing residential parcels and drive routes used for panoramic data collection.
During each field visit, we collected street‐level, 360° imagery using a GoPro Max panoramic camera mounted on a rigid roof–rack mast (approximately 2 m above ground level) on a passenger vehicle. Videos were recorded at the camera’s native panoramic resolution (5.6K) at 30 fps to support frame‐level alignment with GPS. Each recording was accompanied by a synchronized per‐video GPS log (CSV) capturing latitude/longitude, speed, and course, enabling subsequent geospatial matching of frames to reference parcels (Section III-B). The camera was leveled and oriented with the vehicle’s leftward heading to better capture the building facades, with horizon lock enabled to stabilize the panoramic stream. For each campaign, operators maintained a simple time‐stamped field log noting start/stop points, detours, obstructions, and environmental conditions to facilitate post hoc quality control and alignment. All data were captured from public rights‐of‐way; no entry onto private property occurred. To protect privacy, personally identifying details (e.g., faces and license plates) were blurred or removed prior to the dissemination of imagery.
III-B Data pre-processing
This approach assembles a street-level video and GPS dataset to generate parcel-linked facade views suitable for downstream vision-language prompting and temporal inference. The workflow includes matching reference buildings to vehicle GPS tracks at frame resolution, extracting the corresponding video frames, estimating vehicle heading, and dewarping panoramic images to centered rectilinear views.
Reference points are converted to a projected CRS for metric operations (EPSG:5070) and buffered by 25 m to form search regions. For each GPS file, latitude/longitude samples are transformed to the same CRS and converted to point geometries. An R-tree spatial index over buffered reference regions retrieves candidate buildings per GPS point; among candidates, the nearest reference by Euclidean distance is kept. For each match, we record the reference index, the video identifier, the vehicle coordinates from the GPS sample, and a frame index defined as the row position in the GPS file (frame number). Matched rows are merged back to the reference attributes (reprojected to EPSG:4326), duplicates are removed, a unique ObjectId (1..nn) is assigned, and the video name is cleaned to a base identifier. The result is saved as a table for frame extraction.
III-B1 View rectification
Figure 2: Visualization of the view rectification process. (a) The top-down view illustrates the geometric parameters used for rectification, including the vehicle’s position $(x_{1},y_{1})$ and heading $(\theta_{1})$, and the building’s centroid $(x_{2},y_{2})$. The bearing to the building $(\theta_{2})$ is calculated, and the yaw angle $\alpha$ orients the rectified field of view (green) toward the building. (b) A raw 360-degree panoramic image is processed to generate a rectified, planar view of the building facade, which is then used for analysis.
Frame extraction groups videos by frame number, finding the corresponding media file by trying common extensions and name variants, and obtaining duration and frame rate. Requested frame indices are validated against the total frame count; invalid requests are dropped with reasons recorded. Each valid frame number is converted to a timestamp (frame number/fps) and extracted with FFmpeg as a single high-quality JPEG Frames are named by ObjectId and written to a frames directory.
Since the vehicle’s heading information is unavailable, for each matched record, the process reopens the corresponding GPS track and approximates the vehicle heading at the target frame using nearby samples: averaging a small window of points before and after the frame (for example, $\pm$ 15 frames, approximately 0.5 s at 30 fps), then computes the bearing from the mean “before” position to the mean “after” position. This bearing (degrees) is stored as the vehicle’s current heading. The enriched table is saved for subsequent dewarping.
Using building and vehicle coordinates, the algorithm computes the bearing from the camera to the building and derives a yaw to center the facade: $\text{yaw }\alpha=(\text{bearing to building }\theta_{2})-(\text{vehicle heading }\theta_{1})$ plus a camera-specific calibration offset ($\theta_{\text{calib}}$, $90^{\circ}$ in our runs). Yaw is normalized to [180,180][-180,180] degrees. For panoramic frames (wide aspect ratio indicating equirectangular input), the system dewarps to a rectilinear view using FFmpeg’s v360 filter with fixed parameters (horizontal field of view = 90, aspect ratio = 16:9, output width = 1920 px). Non-panoramic frames are passed through unchanged. Outputs are written to a separate directory for rectified frames, and existing outputs may be skipped to allow incremental runs. Figure 2 illustrates the geometric relationships and rectification process.
Quality control tracks invalid frame indices, missing GPS rows, and dewarping skips, and records whether a frame was rectified or copied. The process yields a raw frame and a rectified frame centered on the target facade, together with a table containing ObjectId, reference attributes, video name, frame number, vehicle coordinates, and estimated heading per matched building-view.
The data processing pipeline includes several key calculations. The bearing calculation between two geographic points $(lat_{1},lon_{1})$ and $(lat_{2},lon_{2})$ follows the standard great-circle bearing formula:
$\Delta lon$ $=lon_{2}-lon_{1}$ (1)
$y$ $=\sin(\Delta lon)\cdot\cos(lat_{2})$ (2)
$x$ $=\cos(lat_{1})\cdot\sin(lat_{2})$
$\quad-\sin(lat_{1})\cdot\cos(lat_{2})\cdot\cos(\Delta lon)$ (3)
$\theta$ $=\text{atan2}(y,x)\cdot\frac{180}{\pi}$ (4)
where $\theta$ is normalized to $[0,360]$ degrees. For spatial matching, reference buildings are buffered by 25 m in projected coordinates (EPSG:5070), and an R-tree spatial index enables efficient retrieval of candidate matches within buffer bounds. Vehicle heading estimation averages GPS positions within a temporal window of $\pm 15$ frames ($\approx 0.5$s at 30 fps) and applies the bearing formula to the mean positions. Frame number $f$ is converted to timestamp $t$ via $t=f/\text{fps}$, where fps is the video frame rate. For view rectification, the yaw angle $\alpha$ that centers the building in the rectified view is computed as:
$\alpha=(\theta_{2}-\theta_{1}+\theta_{\text{calib}})\bmod 360$ (5)
where $\theta_{2}$ is the bearing from camera to building, $\theta_{1}$ is the vehicle heading, and $\theta_{\text{calib}}=-90^{\circ}$ is the camera calibration offset. The yaw is normalized to [180,180][-180,180] degrees before application to the FFmpeg v360 filter.
The resulting rectified facade views constitute the inputs to the VLM prompting module (Section III-C) and the temporal comparison and recovery mapping (Section III-D). Calibration offset, field-of-view, and window size are tunable; reported analyses use values set a priori for the study area and collection period.
To evaluate the performance of the method, rectified facade images were independently labeled by two annotators using a three-class system (Occupied, Not Occupied, Uncertain) guided by a written code book defining observable cues (e.g., debris piles, access obstructions, emergency markings, active repairs, vehicle presence). The GoPro operator reviewed all of the labels against field logs and raw panoramas to verify context and visibility and made the final determination of the occupancy status for each frame. Frames labeled as Uncertain or with insufficient visibility were excluded from primary accuracy metrics. For locations visited in both campaigns, labels were assigned per campaign independently to support temporal comparisons.
III-C VLM prompting strategy
The proposed building occupancy assessment in this work employs vision-language models to analyze the rectified facade images from the pre-processing pipeline. Two distinct strategies are implemented and compared: a single-stage vision-only approach and a two-stage vision-reasoning pipeline (Figure 3).
Figure 3: Overview of the two prompting strategies for building occupancy classification. (a) Single-stage baseline: a vision-language model (VLM) extracts nine visual attributes, and a deterministic scoring rule with threshold $\tau$ produces the final label. (b) Two-stage strategy: the same VLM extracts the attributes; a text-only reasoning LLM applies explicit rules and few-shot exemplars to generate a conservative final decision.
The first strategy, which serves as a baseline, applies a fixed-count rule over eight risk indicators (all schema fields except vehicle_presence). Let $r$ denote the number of risk indicators that are true, and let v=1v{=}1 if a curbside vehicle is visible (vehicle_presence=true) and 0 otherwise. The decision policy implemented in code is equivalent to
$\text{predict Not Occupied}\iff r-v\geq\tau,\quad\text{with }\tau=2.$ (6)
Thus, the effective threshold is two when no vehicle is visible and three when a vehicle is present (i.e., one point is subtracted for visible vehicles to reflect the prior that cars correlate with occupancy). Table I summarizes the rules used in all experiments.
TABLE I: Baseline threshold ($\tau$) with vehicle adjustment. Risk indicators are: house_destruction, structural_damage, exterior_debris, open_doors_windows, site_accessible=false, exterior_mud, emergency_markings, major_repairs.
Vehicle visible? Effective $\tau$ (risk indicators needed) Decision summary
No (vehicle_presence=false) $\tau=2$ Not Occupied if $r\geq 2$; else Occupied
Yes (vehicle_presence=true) $\tau=3$ Not Occupied if $r\geq 3$; else Occupied
The choice of $\tau{=}2$ reflects a conservative rule: isolated single cues (e.g., a small debris pile) are insufficient for a Not Occupied decision without corroboration, and visible vehicles raise the required count by one to reduce false positives in occupied but repairing homes. This keeps the baseline transparent and tunable; future work will sweep the threshold to empirically validate the operating point.
The second strategy introduces a separate decision-making stage. In the feature-extraction stage, the same multi-modal model analyzes each rectified image with a prompt that requests only observable visual evidence (the nine attributes above), without inferring occupancy. In the subsequent decision stage, we use the language model (GPT-4o) in text-only mode to map the extracted attributes to a final occupancy label via explicit reasoning rules and few-shot learning examples (in the Appendix section). These rules and exemplars (nine cases covering diverse damage and occupancy scenarios) encode domain expertise. The decision policy is intentionally conservative, favoring “Not Occupied” when evidence is ambiguous or suggests potential habitability concerns. This two-stage approach is referred to as the “Two-stage” strategy in the following analysis.
This two-stage approach offers several methodological advantages over direct classification. The separation of perception and reasoning enables systematic evaluation of visual feature extraction accuracy independent of decision logic. The structured attribute representation provides interpretable intermediate outputs that facilitate error analysis and model debugging. The explicit reasoning stage allows incorporation of domain knowledge and conservative decision policies without requiring retraining of the vision components. Additionally, the standardized attribute schema enables consistent evaluation across different building types and damage patterns. The attributes used in the VLM extraction stage and their typical visual cues are summarized in Table in the Appendix section. Note that Emergency responders often spray-paint a large “X” on building exteriors during search-and-rescue operations, indicating that the structure was inspected and often that it was damaged or uninhabitable.
Both strategies are evaluated on the same dataset of rectified building facade images to assess their relative performance in post-disaster occupancy assessment tasks.
All language-model calls used the OpenAI API via a chat/responses endpoint that accepts role-based messages and images. For fairness, both strategies shared an identical vision–attribute extraction stage (same multimodal model, standard schema, and prompt); only the decision stages are different. The vision extractor was invoked with a system message and a user message (image + standard schema) at temperature=0, max_tokens=200; the decision stage used a text-only call at temperature=0, max_tokens=5. No explicit response_format (JSON mode) was requested—JSON structure was enforced by instruction. The one-stage baseline applied a deterministic scoring rule to the extracted attributes. The two-stage method used a text-only LLM decision with the system message “You are a helpful assistant that analyzes images and outputs structured JSON.”, $\texttt{temperature}=0$, max_tokens=5, and the few-shot exemplars in the Appendix section. No seed was set, and other sampling parameters used defaults (e.g., top_p). The code does not implement automatic JSON validation or retry-on-failure; responses were consumed as returned.
III-D Temporal comparison and recovery mapping
This stage derives parcel‑level occupancy change descriptors between the two field campaigns for each inference strategy and prepares data structures for recovery mapping. Inputs are per‑parcel visit labels from the VLM prompting stage (Section III-C) for both the single‑stage (vision scoring) and two‑stage (vision + reasoning) strategies, together with human ground‑truth annotations for the subset of parcels manually labeled in each campaign. After visibility and quality filtering, parcel identifiers are joined across visits in chronological order (V1 for Visit 1 and V2 for Visit 2). When multiple rectified frames exist for a parcel within a visit, each strategy first produces frame‑level labels; a visit‑level label is then assigned by majority vote. Ties, mutually contradictory evidence, or any frame flagged Uncertain produce a conservative Not Occupied assignment; if all usable frames are Uncertain, the visit label is set to Uncertain and excluded from primary accuracy and change metrics. The same consolidation rule is applied to ground‑truth frame annotations.
For parcels with valid (Occupied or Not Occupied) labels at both visits, we define a change class given by the ordered pair (V1, V2). The four resulting classes are: Recovered (from Not Occupied to Occupied), Deteriorated (from Occupied to Not Occupied), Stable–Occupied (stay as Occupied), and Stable–Not Occupied (stay as Not Occupied), summarized in Table II. Parcels with an Uncertain or missing label at either visit are excluded from change summaries for that strategy but still counted in coverage statistics (fraction of parcels yielding a usable change).
TABLE II: Parcel occupancy change classes between Visit 1 (V1) and Visit 2 (V2).
Change Definition (V1 to V2)
Recovered Not Occupied to Occupied
Deteriorated Occupied to Not Occupied
Stable–Occupied Occupied to Occupied
Stable–Not Occupied Not Occupied to Not Occupied
Structured outputs of this stage are: (1) a visit‑level label table containing, for each parcel and strategy (and ground truth where available), the consolidated V1 and V2 labels with exclusion flags; and (2) a parcel change table enumerating change classes for parcels with valid labels at both visits. These outputs serve as inputs to the evaluation metrics and numerical results in Section IV.
III-E Model Performance
We report 95% confidence intervals (CIs) for Precision, Recall, F1, and Cohen’s $\kappa$ via paired bootstrap resampling at the parcel level. We emphasize Recall and F1 because the dataset is moderately imbalanced between Occupied and Not Occupied, making these metrics more informative than raw accuracy. Cohen’s $\kappa$ complements them by accounting for agreement beyond chance, which is important in imbalanced settings. Between-method differences are defined as $\Delta=\text{two-stage}-\text{one-stage}$ and are assessed with paired bootstrap tests (reporting $\Delta$ and 95% CIs) and McNemar’s test applied separately to Visit 1 and Visit 2 contingency tables. We interpret $p$-values as evidence for or against a difference in error rates, while effect sizes and their CIs guide practical conclusions.
IV Results
IV-A VLM prompting pipeline performance
The original dataset included 578 building facades from Visit 1 and 621 from Visit 2. After filtering out images labeled as “Unknown” due to obstruction, blur, poor viewing angle, or no building present, the usable sample comprised 471 facades for Visit 1 and 443 for Visit 2. These filtered counts define the set of parcels used for model evaluation across both strategies. Model performance is compared to ground truth in several ways. The Ground Truth, predictions by the Two-stage strategy, and by the One-stage strategy are shown in Figure 4. The figure shows the occupancy status for each parcel at both visits. Ground truth vacancies declined from 85 to 77 (–8; –9.4%), indicating a modest decrease in vacancies between the two field surveys. The Two-stage strategy shows the same trend but with greater magnitude (75 to 63; –12; –16.0%), while the One-stage strategy shows an even larger drop (70 to 54; –16; –22.9%). However, both models overestimate occupancy relative to ground truth at each visit (Two-stage: –10 V1 / –14 V2; One-stage: –15 V1 / –23 V2), with larger overestimates at V2 consistent with under-predicted vacancy. Total counted parcels differ between visits because parcels with Uncertain or missing labels are excluded, but the relative shifts still show an occupancy contraction overstated by the One-stage baseline. These distributional biases signal the net recovery bias and change agreement results reported below.
Figure 4: Occupancy status overview for Ground Truth, Two-stage predictions, and One-stage predictions across both visits.
Figure 5: Confusion matrices for both visits, comparing model predictions to ground truth. Not Occupied is treated as a positive class.
For each strategy, we compute per‑visit confusion matrices (shown in Figure 5; Not Occupied is treated as positive class) and report accuracy, precision, recall, F1, and Cohen’s $\kappa$ in Table IV. Table IV summarizes performance with 95% CIs, and Table III reports paired significance tests.
TABLE III: Statistical significance tests for method comparisons. McNemar tests compare per-visit error patterns; paired bootstrap tests compare metric differences.
Test Statistic p-value Interpretation
McNemar (Visit 1) 0.571 0.450 Not significant
McNemar (Visit 2) 0.000 1.000 Not significant
Paired Bootstrap ($\Delta F_{1}$) +0.026 0.188 Not significant
Paired Bootstrap ($\Delta\kappa$) +0.029 0.226 Not significant
TABLE IV: Method performance comparison with bootstrap confidence intervals. Not Occupied is treated as the positive class.
Metric One-stage One-stage 95% CI Two-stage Two-stage 95% CI Difference
Precision 0.943 [0.886, 1.000] 0.927 [0.873, 0.998] -0.016
Recall 0.728 [0.648, 0.862] 0.781 [0.707, 0.902] +0.053
F1-Score 0.822 [0.759, 0.900] 0.848 [0.795, 0.930] +0.026
Cohen’s $\kappa$ 0.789 [0.718, 0.884] 0.818 [0.756, 0.916] +0.029
As shown in Table IV, the two-stage strategy exhibits a modest precision-recall trade-off relative to the one-stage baseline. It achieves higher recall at the cost of slightly lower precision, resulting in numerically higher F1 and Cohen’s $\kappa$ scores. However, as shown in Table III, these differences are not statistically significant. Both McNemar’s tests on per-visit error patterns and paired bootstrap tests on the performance metrics yield non-significant p-values, with confidence intervals for the differences spanning zero. This suggests that while the two-stage method yields slightly higher recall and agreement, the two strategies are statistically indistinguishable at the current operating point.
Figure 6: Per‑visit overall accuracy: (a) One-stage strategy (95.5% V1, 92.1% V2); (b) Two-stage strategy (96.2% V1, 93.2% V2).
Overall accuracies (Figure 6) decline modestly from V1 to V2 for both strategies (Two-stage -3.0 pp (percentage point) vs One-stage -3.4 pp), with the Two-stage maintaining a consistent  0.7–1.1 pp advantage each visit. The parallel drop suggests a shared increase in classification difficulty (e.g., more subtle post‑repair cues) rather than strategy‑specific drift.
Direct comparison of the two strategies focuses on expected advantages of the two‑stage (vision + reasoning) method: fewer false Not Occupied labels at a single visit (improved Occupied recall), higher precision for Deteriorated changes by propagating fewer spurious vacancies, improved precision for Recovered (fewer artificial recoveries caused by an incorrect Not Occupied baseline at V1), and higher overall change $\kappa$. Disagreements between strategies are flagged for manual review and spatially highlighted.
(a) Case 1: both strategies correct (Not occupied).
(b) Case 2: both strategies correct (Occupied).
(c) Case 3: disagreement between strategies.
(d) Case 4: both strategies misclassified.
Figure 7: Examples of rectified building facades with ground truth and model predictions. Sub-figures illustrate a range of outcomes: (a–b) correct predictions, (c) strategy disagreement, and (d) residual error, which is probably caused by mistakenly recognizing the logo on the building as “X” marks. Only attributes that are true are shown above each figure.
Figure 8: Error pattern analysis comparing correctness across parcels: Both Correct, Both Wrong, only One-stage wrong, or only Two-stage wrong (counts with percentages).
Figure 7 shows four representative cases. In panel (a), both strategies correctly classify the parcel as Not Occupied, supported by clear exterior debris and boarded openings. In panel (b), both strategies correctly label an intact home as Occupied. Panel (c) illustrates a disagreement: the one-stage baseline predicts Occupied while the two-stage reasoning strategy assigns Not Occupied, reflecting its more conservative decision rule. Finally, panel (d) shows a residual error where both strategies misclassify due to confounding visual cues (in this case, the model mistakenly interpreted the logo on the door as ”X” markers). Together, these examples highlight how the methods align with ground truth in most straightforward cases, but also where conservative reasoning can alter outcomes or where both models still struggle.
Figure 8 shows that most parcels (211; 87.9%) are correctly classified by both strategies. Joint errors are more frequent (18; 7.5%), and asymmetric errors slightly favor the two-stage strategy (one-stage only wrong: 7; 2.9% vs two-stage only wrong: 4; 1.7%). This pattern corroborates the two-stage model’s higher recall and agreement, while highlighting that residual challenges concentrate in a small set of ambiguous parcels (Both Wrong) suitable for targeted prompt refinement or data augmentation.
IV-B Occupancy change analysis
We analyze parcel‑level occupancy change between V1 and V2 using the four classes defined in Table II (Stable Occupied, Stable Not Occupied, Recovered, Deteriorated). Figure 9 (a)–(c) show Recovered / Deteriorated counts: ground truth 20 / 6 (net +14), Two-stage 23 / 9 (net +14; matches net recovery), One-stage 24 / 6 (net +18; overstates recovery by +4). Thus, both strategies capture the direction of net improvement; the One-stage inflates recoveries, while the Two-stage balances additional true recoveries with extra deteriorations.
Figure 9 (d) partitions parcel outcomes: Perfect Agreement: No Change (197; 82.1%), Perfect Agreement: Change (17; 7.1%), partial agreement where only one model matches ground truth (Two-stage only 8; 3.3%, One-stage only 4; 1.7%), and both models agree but wrong (14; 5.8%). The last two groups can be a guide for the targeted review.
Figure 9: Change analysis. (a) Ground‑truth Recovered vs. Deteriorated counts (n=26). (b) One-stage predicted changes (n=30). (c) Two-stage predicted changes (n=32).
TABLE V: Parcel-level occupancy change agreement categories (confident parcels with both visits).
Category Count Rate (%)
Perfect Agreement: No Change 197 82.1
Perfect Agreement: Change 17 7.1
GT + Two-stage Agree 8 3.3
GT + One-stage Agree 4 1.7
Methods Agree, GT Differs 14 5.8
Total 240 100.0
Spatial heterogeneity and remaining uncertainty are visualized in Figure 10, which maps (a) Visit 1 ground‑truth occupancy, (b) Visit 2 ground‑truth occupancy, (c) ground‑truth change classes (Recovered, Deteriorated, Unchanged), and (d) spatial distribution of Two-stage prediction accuracy (both visits correct, one visit correct, both wrong). This multi‑panel view highlights localized pockets of recovery and deterioration and reveals where residual model errors cluster geographically, indicating priority areas for data or prompt refinement.
Figure 10: Geographic distribution of occupancy and model accuracy: (a) Visit 1 ground‑truth status; (b) Visit 2 ground‑truth status; (c) ground‑truth change classes (Recovered, Deteriorated, Unchanged); (d) Two-stage prediction accuracy by parcel (both visits correct, one visit correct, both wrong).
Together, these analyses contextualize raw change counts with spatial concentration, statistical confidence, and directional bias diagnostics, informing both methodological refinement and operational decision making.
We assessed spatial autocorrelation with the global Moran’s $I$ for four spatial fields: parcel occupancy at Visit 1 and Visit 2, the ground-truth change field, and a binary accuracy indicator (1 = correct, 0 = incorrect). Positive $I$ with small $p$-values indicates spatial clustering rather than spatial randomness. Table VI summarizes the results.
TABLE VI: Global Moran’s II statistics for parcel-level layers. Positive values indicate spatial clustering. The expected value under the randomization null is E[I]=1/(n1)E[I]=-1/(n-1); here $E[I]\approx-0.0042$ because $n=240$.
Layer Moran’s II Expected II $p$-value
Visit 1 occupancy 0.3704 0.0042-0.0042 0.0000
Visit 2 occupancy 0.2834 0.0042-0.0042 0.0000
Ground truth changes 0.0799 0.0042-0.0042 0.0048
Prediction accuracy 0.0600 0.0042-0.0042 0.0313
Under the standard randomization null, the expected global Moran’s $I$ equals 1/(n1)-1/(n-1), where $n$ is the number of parcels in the analysis; using our common parcel mask ($n=240$) yields $E[I]\approx-0.0042$ for each layer. While the associated $p$-values confirm that clustering departs from spatial randomness, we interpret $I$ primarily as an effect size and rely on local statistics and maps to indicate where clustering concentrates.
These statistics indicate significant clustering for all four layers. Clustering is strongest for occupancy states (Visit 1 I=0.3704I{=}0.3704, Visit 2 I=0.2834I{=}0.2834) and weaker, but still significant, for the change field (I=0.0799I{=}0.0799) and the accuracy surface (I=0.0600I{=}0.0600). Practically, residual errors are not uniformly scattered; they concentrate in localized pockets, which supports prioritizing targeted audits of those neighborhoods.
V Discussion
Aggregate comparisons with uncertainty and paired tests indicate that, at our chosen operating point, the two-stage strategy achieves higher recall and slightly higher F1 and Cohen’s $\kappa$ than the one-stage baseline, at the cost of a small reduction in precision. These differences are not statistically significant (McNemar and paired bootstrap), so we regard the two strategies as comparable in overall agreement. The operational value of the two-stage design lies in interpretability and controllability: exposing intermediate reasoning and attribute signals supports risk-aware thresholding and targeted review of spatial error clusters. While our method measures parcel-level occupancy change, these shifts provide a practical proxy for community recovery when interpreted alongside other indicators.
Beyond accuracy, the approach reduces survey time and labor. Across two campaigns we produced usable facades for 914 residential parcels at roughly 90 parcels per survey hour. It is an order-of-magnitude gain over house-to-house assessments, which reinforces its value as a scalable complement to manual surveys when rapid neighborhood information is required.
Accuracy declined from Visit 1 to Visit 2 for both methods, likely due to subtler post-repair cues and weaker signals. Few parcels were uniquely misclassified by the reasoning approach, suggesting the added layer increased signal rather than noise. Extending to longer-term recovery monitoring will require richer temporal context beyond facade imagery alone.
Change detection highlights the benefits of reasoning. Both approaches captured the direction of occupancy, but the baseline exaggerated improvement by propagating an inflated vacancy estimate. The reasoning strategy balanced additional true recoveries with a few deteriorations, producing a more faithful estimate of net change—important for field tasking and resource planning.
Spatial mapping adds interpretability by revealing residual errors clustered in small pockets, while large areas show consistent agreement with ground truth. These patterns likely reflect local building or damage typologies and enable efficient triage by directing review to high‑uncertainty regions. The pipeline scales raw street‑view video into parcel‑level states, limiting human oversight to ambiguous cases, and its modular prompts and model weights support adaptation as models improve.
Limitations include evaluation in a single event area and exclusion of uncertain or missing labels. Only two visits constrain temporal inference and may miss nonlinear recovery (e.g., temporary re‑vacancy), and environmental variation (illumination, foliage, seasonality) was not modeled, limiting transferability over time. Reasoning traces were not systematically audited, and both strategies share the same VLM backbone, so results reflect prompt and architectural choices rather than new model training.
These limitations suggest next steps: (i) additional visits and simple temporal models; (ii) using spatial disagreement maps to target data collection; (iii) prompt and attribute refinements for recurring failure modes (e.g., boarded or tinted windows); (iv) fusing VLM outputs with GIS/building attributes and calibrated probabilities for risk-aware thresholds; and (v) auditing reasoning traces and codifying privacy policies.
In summary, inserting a concise reasoning stage yields more balanced and directionally accurate monitoring of post-disaster vacancy than a single-stage baseline. Improvements concentrate on vacancy recall and net-recovery estimation, while residual errors remain geographically localized and thus tractable for human review. Expanding geographic scope, adding temporal depth and uncertainty, and stress-testing environmental robustness are key steps toward operational deployment.
VI Conclusion
Automated occupancy monitoring is a crucial step toward rapid recovery tracking. Leveraging street-view imagery (SVI) makes this operational because it observes the facade and access cues—blocked entries, mud lines, emergency placards, active repairs—that nadir imagery cannot see. We presented a street-level, parcel-scale pipeline that pairs panoramic rectification with vision–language prompting to assess post-disaster building occupancy and track change between visits, comparing a transparent one-stage rule with a two-stage design that separates perception from conservative reasoning.
Across two visits, both strategies achieved high agreement with ground truth. The two-stage method traded a slight precision drop for higher recall and reproduced the ground-truth net recovery (+14), while the one-stage baseline overstated improvement. Spatial diagnostics (global Moran’s $I$ and error maps) revealed significant clustering of both occupancy states and residual errors, enabling targeted review of localized pockets.
Operationally, separating perception and reasoning surfaces interpretable intermediate signals and supports risk-aware thresholding and human-in-the-loop oversight. Limitations include a single event area, two field visits, exclusion of uncertain labels, unmodeled environmental variation, and reliance on one VLM backbone. Next steps include expanding geographic coverage, adding temporal depth and calibrated probabilities, auditing reasoning traces, and fusing VLM outputs with lightweight GIS and building attributes.
In sum, the approach provides a scalable way to convert raw street-view video into parcel-level occupancy states and recovery indicators. While aggregate differences between strategies are not statistically significant, the two-stage design offers practical advantages in interpretability and controllability, and the spatial analyses identify where to focus limited field and labeling resources for reliable, operational deployment.
Acknowledgments
Funding
This material is supported by the National Science Foundation under the CAREER grant (No. 1846069) and the North Carolina General Assembly through the North Carolina Policy Collaboratory.
Data and Code availability
The data that support the findings of this study are publicly available at Huggingface, and the code is available at GitHub.
Appendix A VLM Prompts and Few-shot Examples
Listing 1: Actual vision-only prompt used (attribute extraction)
Analyze the image and answer with a JSON object using EXACTLY these keys and boolean values (true/false). Do not add, remove, or rename keys. Return only the JSON object, no prose.
house_destruction”: true/false,
structural_damage”: true/false,
exterior_debris”: true/false,
open_doors_windows”: true/false,
site_accessible”: true/false,
exterior_mud”: true/false,
emergency_markings”: true/false,
major_repairs”: true/false,
vehicle_presence”: true/false
Listing 2: Expected JSON string format (canonical schema)
{”house_destruction”: false,
structural_damage”: true,
exterior_debris”: true,
open_doors_windows”: false,
site_accessible”: true,
exterior_mud”: false,
emergency_markings”: false,
major_repairs”: false,
vehicle_presence”: true}
A-A Prompts for decision stage (few-shot)
Listing 3: Decision-stage prompt with few-shot examples (using canonical schema)
You are an expert in post-disaster building occupancy assessment. Given a buildings attributes in JSON (keys listed below), decide if it is Occupied or Not Occupied’.
Consider all evidence: for example, some exterior mud may coexist with occupancy; parked vehicles can indicate occupancy when other signs are mixed; extensive roof or wall repairs may indicate temporary non-occupancy. If the evidence is mixed or unclear, prefer Not Occupied’. Be conservative: if there is significant indication that a building might not be occupied (e.g., destruction, inaccessibility, visible abandonment, or multiple risk factors), classify it as Not Occupied’. Classify as Occupied only if the building appears livable and there are no clear signs of uninhabitability.
The JSON fields you will receive:
house_destruction”: bool,
structural_damage”: bool,
exterior_debris”: bool,
open_doors_windows”: bool,
site_accessible”: bool,
exterior_mud”: bool,
emergency_markings”: bool,
major_repairs”: bool,
vehicle_presence”: bool
Here are some examples:
Example 1:
house_destruction”: false,
structural_damage”: false,
exterior_debris”: false,
open_doors_windows”: false,
site_accessible”: true,
exterior_mud”: false,
emergency_markings”: false,
major_repairs”: false,
vehicle_presence”: true
Occupied
Example 2:
house_destruction”: true,
structural_damage”: true,
exterior_debris”: true,
open_doors_windows”: false,
site_accessible”: true,
exterior_mud”: false,
emergency_markings”: false,
major_repairs”: false,
vehicle_presence”: false
Not Occupied
Example 3:
house_destruction”: false,
structural_damage”: true,
exterior_debris”: true,
open_doors_windows”: true,
site_accessible”: true,
exterior_mud”: false,
emergency_markings”: false,
major_repairs”: false,
vehicle_presence”: true
Not Occupied
Example 4:
house_destruction”: false,
structural_damage”: true,
exterior_debris”: false,
open_doors_windows”: false,
site_accessible”: true,
exterior_mud”: false,
emergency_markings”: false,
major_repairs”: true,
vehicle_presence”: false
Not Occupied
Example 5:
house_destruction”: false,
structural_damage”: false,
exterior_debris”: false,
open_doors_windows”: false,
site_accessible”: true,
exterior_mud”: false,
emergency_markings”: true,
major_repairs”: false,
vehicle_presence”: false
Not Occupied
Example 6:
house_destruction”: false,
structural_damage”: false,
exterior_debris”: false,
open_doors_windows”: false,
site_accessible”: true,
exterior_mud”: false,
emergency_markings”: false,
major_repairs”: false,
vehicle_presence”: false
Occupied
Example 7:
house_destruction”: true,
structural_damage”: false,
exterior_debris”: false,
open_doors_windows”: false,
site_accessible”: true,
exterior_mud”: false,
emergency_markings”: false,
major_repairs”: false,
vehicle_presence”: false
Not Occupied
Example 8:
house_destruction”: false,
structural_damage”: false,
exterior_debris”: false,
open_doors_windows”: false,
site_accessible”: false,
exterior_mud”: false,
emergency_markings”: false,
major_repairs”: false,
vehicle_presence”: false
Not Occupied
Example 9:
house_destruction”: false,
structural_damage”: false,
exterior_debris”: false,
open_doors_windows”: false,
site_accessible”: true,
exterior_mud”: true,
emergency_markings”: false,
major_repairs”: false,
vehicle_presence”: true
Occupied
Now, decide for this building:
{…JSON from vision model…}
Output only one token: Occupied or Not Occupied’.
Output only one token: Occupied or Not Occupied’.
TABLE VII: Attribute schema used in VLM extraction: definitions and typical observable cues.
Attribute in prompts Typical visual cues (examples)
house_destruction Severe or total loss of habitability: collapse, missing major wall sections, roof largely gone, house off foundation.
structural_damage Visible structural or envelope damage short of destruction: partial roof damage, compromised porch, broken/boarded windows, leaning walls, major siding loss.
exterior_debris Disaster or cleanup debris in yard/drive/curb: piles of damaged contents, construction rubble, spoiled furniture, scattered materials (exclude small ordinary clutter).
open_doors_windows Door or window visibly open (ajar) or missing, suggesting exposure; excludes normal closed / reflective windows where interior not visible.
site_accessible Parcel frontage unobstructed (TRUE) vs blocked (FALSE) by fencing, gate, caution tape, barricade, large dumpster/trailer preventing normal entry.
exterior_mud Distinct flood or sediment deposits: mud lines on siding, dried silt on walls/steps/yard, uniform brown coating inconsistent with normal dirt.
emergency_markings Official post-disaster markings: spray-painted search/inspection tags (e.g., X marks), placards, colored notices on doors/windows; exclude ordinary numbers or graffiti.
major_repairs Active or recent substantial repair work: tarps on roof, exposed framing, scaffolding, large construction materials (lumber stacks, shingles bundles), contractor equipment.
vehicle_presence One or more vehicles visibly parked in driveway, curb, or immediate lot area (cars, trucks, vans). Construction-only equipment without a personal vehicle may be excluded.